# Exploring the 6G Potentials: Immersive, Hyper Reliable, and Low-Latency Communication

### v2

@@ I Introduction
Recently, the fifth-generation (5G) networks have become operational in many countries worldwide, while researchers are working on studies related to the sixth-generation (6G) wireless networks. These studies are inspired by potential use cases, requirements, and unresolved challenges. The previous generations of wireless networks have already provided significant enhancements that enabled a vast range of novel applications compared to those supported by legacy networks. Nevertheless, according to the International Telecommunication Union’s (ITU) vision for international mobile telecommunications 2030 (IMT-2030), 6G usage scenarios are envisioned as enhanced versions of existing 5G scenarios, alongside entirely new use cases that support the deployment of innovative applications [1]. Additionally, 6G wireless networks are expected to cater to future use cases and applications that require the concurrent fulfillment of high mobility, immersive data rates, hyper-reliability, and low-latency. This necessitates the introduction of significantly more advanced network architectures and enabling technologies. We refer to this type of service class as immersive, hyper reliable, and low-latency communication (IHRLLC), which is also the focus of this article.
@@
  • We present a numerical simulation-based experiment to validate the high throughout via ultra-massive multiple-input multiple-output (umMIMO)-aided THz communication.
@@ III Envisioned Enabling Technologies for IHRLLC
Figure 2: Per user throughput and BER versus the number of users for umMIMO-enabled THz communication.
@@
Higher frequency bands, such as mmWave and THz, enable the use of umMIMO antenna designs that can accommodate a significantly greater number of antenna elements, compared to the traditional MIMO architectures. As a result of implementing the umMIMO technology and enhanced beamforming techniques, high-gain narrow antenna beams can be tailored to individual users. These beams help overcome the substantial FSPL and short-range communication in THz communications [8]. Fig. 2 plots the throughput and bit error rate (BER) performance under different numbers of antenna elements versus the number of users in the umMIMO-enabled THz communication system. The network consists of a base station and up to $20$ IHRLLC users in a cell of radius $500$ meters, operating at a center frequency of $850$ GHz with bandwidth of $2$ GHz. The network resources are equally and orthogonally divided among users to eliminate interference. It is observed that the throughput and BER values remain within acceptable ranges, supporting reliable and immersive communication. The performance improves with higher-order umMIMO antennas; for instance, upgrading from a $512\times 512$ antenna array to a $1024\times 1024$ array yields slight throughput improvement. This moderate gain is influenced by the logarithmic nature of the relationship, but further enhancements can be expected with even higher-order umMIMO arrays. Nonetheless, achieving such high data rates and low BER for a large number of users with limited resources highlights the necessity of advanced beamforming algorithms and strong error control coding techniques to optimize performance.
@@ III-B Multi-Connectivity through Non-Terrestrial Networks
In the previous sections, various enabling technologies were discussed to facilitate IHRLLC. The implementation of such sophisticated technologies, as well as the proper design of the corresponding parameters, is a critical challenge. Considering the umMIMO antennas and RISs, with a large number of antenna elements along with the challenges of THz communication, designing the optimum parameters through traditional approaches incurs a high computational cost. To this end, AI algorithms can potentially produce revolutionary changes in future wireless networks and overcome computational challenges. In particular, AI can be integrated with wireless networks as a distinct network function or be embedded within existing functions, depending on network demands. In scenarios where AI needs to handle complex, centralized tasks, AI can be deployed as a dedicated network function, responsible for high-level decision-making, and coordination with other functions to optimize the overall network performance. Alternatively, AI can be integrated as a module within existing network functions to enhance their operations. This modular integration enhances the intelligence and adaptability of current architectures without introducing a completely new function. In the following discussion, we investigate promising novel computational-related approaches, including learn-to-optimize (L2O), generative AI (GenAI), and quantum computing technologies as methods to solve the aforementioned computational challenges in 6G wireless networks.
@@ VI Conclusion
Afsoon Alidadi Shamsabadi [SM] (afsoonalidadishamsa@sce.carleton.ca) is a PhD Candidate in the Non-Terrestrial Networks (NTN) Lab in the Systems and Computer Engineering Department at Carleton University, Canada. Her research focuses on the innovative integration of High Altitude Platform Stations (HAPS) with terrestrial wireless networks.
@@
Halim Yanikomeroglu [F] (halim@sce.carleton.ca) is a Chancellor’s Professor of Systems and Computer Engineering in the Non-Terrestrial Networks Lab, at Carleton University, Canada. His group’s focus is the wireless access architecture for the 2030s and 2040s, and non-terrestrial networks. He is a Fellow of EIC (Engineering Institute of Canada), CAE (Canadian Academy of Engineering), and AAIA (Asia-Pacific Artificial Intelligence Association); and a Distinguished Speaker for IEEE ComSoc and IEEE VTS.

### v3

@@ I Introduction
Recently, the fifth-generation (5G) networks have become operational in many countries worldwide, while researchers are working on studies related to the sixth-generation (6G) wireless networks. These studies are inspired by potential use cases, requirements, and unresolved challenges. The previous generations of wireless networks have already provided significant enhancements that enabled a vast range of novel applications compared to those supported by legacy networks. Nevertheless, according to the International Telecommunication Union’s (ITU) vision for international mobile telecommunications 2030 (IMT-2030), 6G usage scenarios are envisioned as enhanced versions of existing 5G scenarios, alongside entirely new use cases that support the deployment of innovative applications [1]. Additionally, 6G wireless networks are expected to cater to future use cases and applications that require the concurrent fulfillment of high mobility, immersive data rates, hyper-reliability [^1] Reliability, as defined by ITU, is the ability to transmit a predefined amount of data successfully within a specified time and probability., and low-latency. This necessitates the introduction of significantly more advanced network architectures and enabling technologies. We refer to this type of service class as immersive, hyper reliable, and low-latency communication (IHRLLC), which is also the focus of this article.
@@
  • We present a numerical simulation-based experiment to validate the high throughput via ultra-massive multiple-input multiple-output (umMIMO)-aided THz communication.
@@ III Envisioned Enabling Technologies for IHRLLC
Figure 2: Per user throughput and BLER versus the number of users for umMIMO-enabled THz communication.
@@
Higher frequency bands, such as mmWave and THz, enable the use of umMIMO antenna designs that can accommodate a significantly greater number of antenna elements, compared to the traditional MIMO architectures. As a result of implementing the umMIMO technology and enhanced beamforming techniques, high-gain narrow antenna beams can be tailored to individual users. These beams help overcome the substantial FSPL and short-range communication in THz communications [8]. Fig. 2 plots the throughput and block error rate (BLER) performance under different number of antenna elements versus the number of users in the umMIMO-enabled THz communication system. The network consists of a base station and up to $20$ IHRLLC users in a cell of radius $500$ meters, operating at a center frequency of $850$ GHz with bandwidth of $2$ GHz. The network resources are equally and orthogonally divided among users to eliminate interference. It is observed that, with sufficient available resources, the throughput and BLER values will be within acceptable ranges, supporting immersive and reliable communication. The performance improves with higher-order umMIMO antennas; for instance, upgrading from a $512\times 512$ antenna array to a $1024\times 1024$ array yields slight throughput improvement. This moderate gain is influenced by the logarithmic nature of the relationship, but further enhancements can be expected with even higher-order umMIMO arrays. Nonetheless, achieving such high data rates and low BLER for a large number of users with limited resources highlights the necessity of advanced beamforming algorithms and strong error control coding techniques to optimize performance.
@@ III-C AI Enablers: L2O and GenAI
In the previous sections, various enabling technologies were discussed to facilitate IHRLLC. The implementation of such sophisticated technologies, as well as the proper design of the corresponding parameters, is a critical challenge. Considering the umMIMO antennas and RISs, with a large number of antenna elements along with the challenges of THz communication, designing the optimum parameters through traditional approaches incurs a high computational cost. To this end, AI algorithms can potentially produce revolutionary changes in future wireless networks and overcome computational challenges. In particular, AI can be integrated with wireless networks as a distinct network function or be embedded within existing functions, depending on network demands. In scenarios where AI needs to handle complex centralized tasks, AI can be deployed as a dedicated network function, responsible for high-level decision-making, and coordination with other functions to optimize the overall network performance. Alternatively, AI can be integrated as a module within existing network functions to enhance their operations. This modular integration enhances the intelligence and adaptability of current architectures without introducing a completely new function. In the following, we investigate promising novel AI enablers, including learn-to-optimize (L2O) and generative AI (GenAI) as methods to solve the aforementioned computational challenges in 6G wireless networks.
@@ VI Conclusion
Afsoon Alidadi Shamsabadi [SM] (afsoonalidadishamsa@sce.carleton.ca) is a PhD Candidate in the Carleton-NTN (Non-Terrestrial Networks) Lab in the Systems and Computer Engineering Department at Carleton University, Canada. Her research focuses on the non-terrestrial networks and the innovative integration of High Altitude Platform Stations (HAPS) with terrestrial wireless networks.
@@
Halim Yanikomeroglu [F] (halim@sce.carleton.ca) is a Chancellor’s Professor in the Department of Systems and Computer Engineering at Carleton University, Canada, and he is the Director of Carleton-NTN (Non-Terrestrial Networks) Lab. His group’s focus is the wireless access architecture for the 2030s and 2040s, and non-terrestrial networks. He is a Fellow of EIC (Engineering Institute of Canada), CAE (Canadian Academy of Engineering), and AAIA (Asia-Pacific Artificial Intelligence Association).